OpenStudy (anonymous):

Ok, got half this on my own, need help with rest please... The perimeter of a rectangle is 36ft. The length is 4 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting sytem. Let x be the length and y be the width. Write the first equation. 2x + 2y= 36 <=== got that part Write the second equation. x=y+ 4 <==== got that part Need help with these two. What is the length? What is the width?

OpenStudy (anonymous):

replace \(x\) in the first equation by \(y+4\) because the second equation tells you that \(x\) IS \(y+4\) use parentheses carefully, and you get \[2(y+4)+2y=36\]

OpenStudy (anonymous):

now solve for \(y\) in a few steps, first multiply on the left to get \[2y+8+2y=36\] combine terms \[4y+8=36\] and last two steps are to subtract 8 and then divide by 4

OpenStudy (anonymous):

subtract 8 from 36? and then divide that by 4?

OpenStudy (anonymous):

which would be 7?

OpenStudy (anonymous):

then 7+4 gives 11.